# Learning Attentive Neural Processes for Planning with Pushing Actions

### v1

@@ Abstract
Our goal is to enable robots to plan sequences of tabletop actions to push a block with unknown physical properties to a desired goal pose on the table. We approach this problem by learning the constituent models of a Partially-Observable Markov Decision Process (POMDP), where the robot can observe the outcome of a push, but the physical properties of the block that govern the dynamics remain unknown. The pushing problem is a difficult POMDP to solve due to the challenge of state estimation. The physical properties have a nonlinear relationship with the outcomes, requiring computationally expensive methods, such as particle filters, to represent beliefs. Leveraging the Attentive Neural Process architecture, we propose to replace the particle filter with a neural network that learns the inference computation over the physical properties given a history of actions. This Neural Process is integrated into planning as the Neural Process Tree with Double Progressive Widening (NPT-DPW). Simulation results indicate that NPT-DPW generates more effective plans faster than traditional particle filter methods, even in complex pushing scenarios.
@@ I Introduction
In scenarios where robots operate in unstructured, dynamic environments, planning under uncertainty is paramount. Some relevant properties for manipulation, such as object geometry, can be easily observed through sensors. Many other properties, such as inertial properties and friction, require interaction with the object [1]. One way robots can learn these inertial properties is to iteratively interact with objects and observe their behavior.
@@
We focus on an instance of this problem where a robot must compose a sequence of pushes to guide a block from a starting configuration to a goal configuration, without prior knowledge of the center of mass. Instead, it must infer these properties while simultaneously learning a predictive model of the outcomes of pushing actions.
@@
We model our planning challenge as a Partially Observable Markov Decision Process (POMDP), where unobservable physical properties remain constant over time [2]. Many POMDP solutions rely on Bayesian state estimation techniques (e.g., Kalman or particle filters) to update their beliefs as new observations become available. However, the linear assumptions required for Kalman filters do not hold in this scenario, since varying the center of mass can drastically alter rotational outcomes. More general methods, such as particle filters, must be used, despite their high computational cost [3].
@@
To leverage these particle-based belief representations to solve our planning problem, we employ Particle Filter Trees with Double Progressive Widening (PFT-DPW), a Monte Carlo tree search algorithm that performs a randomized search to identify the plan yielding the highest expected reward. In POMDP planning, it is crucial to balance exploration and exploitation [4], because actions that best uncover hidden parameters may not directly move the block toward its goal. PFT-DPW and similar Monte Carlo tree search methods manage this trade-off effectively by systematically exploring uncertain aspects of the system while exploiting known information to achieve high rewards [5].
However, PFT-DPW also has significant limitations due to the use of a particle filter. First, updating the belief scales in cost linearly with the number of particles, which becomes prohibitively expensive when using large neural networks for each update. Since belief updates occur repeatedly as a subroutine within the planning algorithm, these costs accumulate quickly. Second, PFT-DPW is limited in its ability to handle more complex problems: it remains susceptible to the curse of dimensionality, where the required number of particles grows exponentially with the dimension of the underlying distribution, making it difficult to scale to higher-dimensional tasks.
@@
To overcome these problems, we propose a model called the Pushing Neural Process (PNP), which jointly trains an inference network that estimates physical parameters, and a prediction network that learns to output a distribution of push outcomes based on physical properties. The inference network and prediction network are set up in an encoder-decoder architecture, where a representation of the physical properties is learned in the intermediate latent space.
@@
We use our models to derive a new planner, which we call NPT-DPW (Neural Process Tree with Double Progressive Widening). NPT-DPW replaces the slower particle filter used to maintain belief in PFT-DPW with a neural network that consumes the history of a given search and outputs a unified belief distribution. Our results show that, on average, NPT-DPW reaches up to twice as close to the goal location as PFT-DPW and avoids catastrophic failures (such as pushing into prohibited zones) far more often.
@@
Figure 1: Above is a diagram describing the Push Neural Process architecture and its use within NPT-DPW. On the right of the image, the search tree over actions and outcomes is pictured, where at each step an action is selected. We then pass in the history of outcomes ($h$) that we have already observed into the encoder ($q_{\phi}$) of the model to obtain the distribution over the latent parameters, which we sample to get ($z$). We then pass $z$ to the decoder with our new action $a$ to get a distribution over outcomes. We sample from that distribution as an outcome ($o$) for the action $a$. We then add that outcome to the potential outcomes in the search tree.
@@ II Problem Setting
For the objects we aim to push, we categorize their properties into two distinct sets: observable properties, denoted by $x\in\mathcal{X}\subset\mathbb{R}^{n}$, and latent physical properties, denoted by $z\in\mathcal{Z}\subset\mathbb{R}^{m}$, where $m$ is decided empirically based on what results in the strongest model. The observable properties $x$ include features such as the block’s resting orientation, while $z$ captures physical properties such as center of mass. These latent properties are unknown to the robot a priori but play a critical role in determining the outcome of a push. We define a push in terms of an action $a\in\mathcal{A}$, characterized by an approach angle and a push velocity, where the angle is in the range $[0,2\pi]$. Upon completing a push, the robot observes the resulting pose of the block $o\in SE(2)$.
@@ III-A Pushing Neural Process
The encoder consumes $D_{t}$ (of arbitrary length) along with the observable properties $x$ (object geometry) and infers a probability distribution over the latent parameters $z$. To this end, we train a neural network, parameterized by $\phi$, to produce an approximate posterior $q_{\phi}(z|x,D_{t})$ that closely matches the true posterior $p(z|x,D_{t})$.
@@ III-B NPT-DPW
Instead of using a particle filter, our planner, NPT-DPW, uses the Pushing Neural Process to infer the underlying physical parameters and predict outcomes of future actions. To expand the tree, we pass the history of actions and outcomes through the encoder to infer the physical parameters. Actions are sampled as they normally would be in PFT-DPW. Observations are sampled from the distribution computed by the decoder from the inferred parameters and the push under consideration. Thus, every step of the tree search will have runtime complexity of $O(h^{2})$, where $h$ is the size of the history (due to the quadratic complexity of self-attention), rather than $O(n)$, where $n$ is the number of particles used in PFT-DPW. We hypothesize a significant improvement in plan quality within a fixed computational budget, since $h\ll n$ in practice.
@@ IV Experiments
Figure 3: Results plotting average distance from the goal as a function of planning time allocated per action. It demonstrates the effectiveness of the NPT-DPW versus PFT-DPW (with 10/30/100 particles). We conducted trials with 0.5,1,2,3,5, and 10 seconds of time to plan per step.
@@
Figure 4: View of the robot interacting with the block in simulation.
IV-A Scenario
The experiments will be run using three tables with geometry designed to vary the hardness of the pushing problem (Figure 2). Each of these tables defines a different planning problem. The first is a blank-surface table designed to check whether the planner can produce plans that drive the block to the goal. In the second and third scenarios, the movement of the block must be constrained along certain paths to prevent it from falling off the pushing surface. The second one is slightly simpler since it encourages pushing in a straight line, while the third will require the robot to understand how to rotate the block. Solving these instances requires the planner to accurately infer the object’s physical properties. An intelligent planner must sequence information-gathering actions on a wider area before pushing the block along the narrow sections of these tables.
After each push, the robot obtains a reward for the block’s location $(x,y)$ relative to the goal $(g_{x},g_{y})$. We choose the reward function
@@ IV-A Scenario
fallen is a function that returns 1 if the block is off the table, and 0 otherwise. inRange is a function that is $1$ if the block is within 0.2 meters of the goal, and 00 otherwise. We set the discount factor $\gamma=0.6$ to incentivize the planner to prioritize not falling off the table in the next action. An example of a plan created by PFT-DPW is pictured in Figure 5.
@@
We assess performance by fixing the computation time $t$ allowed at each step of the plan. Then for each $t$ in $[0.5,1,2,3,5,10]$ seconds, we ran 15 trials on each map for each planner, where we recorded the distance of the object from the goal after 30 steps. If the object fell off the table, or reached the goal early, we stopped then. We sampled a new center of mass for the block in each trial.
@@
We also varied the number of particles in PFT-DPW, initializing it with $n=10,30$, or $100$, to evaluate NPT-DPW against a range of particle filter–based planners and assess the impact of particle count on performance.
These simulations will be run in the PyBullet simulator (Figure 4), with the robot being programmed to push in a direction while keeping its height and direction consistent throughout the motion. To sample actions for progressive widening, we uniformly sample an angle from $[0,2\pi]$ in world coordinates and push at the velocity of $0.10$ m/sec. Each of the blocks is a rectangular prism with a fixed geometry of 2.5 cm x 2.5 cm x 1.5 cm. The mass of the block is set constant at 0.2 kg. The center of mass of the block is sampled uniformly within the convex hull of the object’s geometry.
Our neural network architecture is based on Attentive Neural Processes [6]. Each push is first encoded by an MLP and processed through multiple layers of self-attention. After that, the outputs of the self-attention layers are aggregated into outputs consisting of a mean and covariance. These represent the mean and diagonal covariance of a multivariate Gaussian over $z$, representing the distribution over latent physical properties. Note that for our problem, the only latent physical properties are the $x$ and $y$ coordinate of the center of mass, but we still set $z\in\mathbb{R}^{5}$ since it empirically led to improved performance. The decoder is structured as multiple linear layers that consume a sample from the latent space and outputs a distribution over possible outcome poses of the block after the push is finished.
@@
Figure 5: This is an example of a plan created by PFT-DPW and run by the simulation on the ”ring” map. Important features that show signs of intelligent planning are the pushes in the beginning in the corner, signifying learning more about the block before trying to move towards the goal. Note the figure has been re-scaled slightly from the real experiments to be more visually clear.
@@ IV-B Results
Our results are shown in Figure 3. When the planner is allocated a second or more to plan, NPT-DPW significantly outperforms PFT-DPW. On the ”beam” map, NPT-DPW performs more than twice as well on average. However, when only half a second per action is allocated, the performance of both models is similarly poor. This is likely because the planners may fail to sample an action angle that moves the block toward the goal.
A plausible explanation for NPT-DPW’s better performance quicker belief updates that allowed for more traversals of the tree. For instance, NPT-DPW always ended up searching at least twice as many actions as PFT-DPW. Additionally, we hypothesize that the self-attention heads in the encoder of the Attentive Neural Process help the model learn structural information about the block. This is because self-attention allows the model to infer relationships between previous push outcomes. In contrast, the particle filter maintains a belief that is updated sequentially with each new observation, which still allows for learning interdependence, but doesn’t directly compare pushes like the neural process does.
We believe the lack of performance differences between various particle counts in PFT-DPW can be attributed to the nature of the problem being solved. In our simulation, small differences in the center of mass are difficult to distinguish, since the push outcome is primarily affected by the relative location of the center of mass to the point of contact. As a result, using only a few particles may yield most of the benefits that additional particles provide, while also allowing the planner to explore more push actions due to cheaper belief updates. This trade-off may explain the trends observed in Figure 3. We expect that with planning times longer than those tested, particle filter–based planners with more particles may eventually surpass smaller models in performance.
Ultimately, our results suggest the viability of using an Attentive Neural Process as an alternative to traditional particle filter–based approaches for planning in POMDPs. In particular, when given less than 10 seconds to plan, NPT-DPW consistently and significantly outperforms PFT-DPW.
@@ V Related Work
Estimating unobserved properties through interaction is a common strategy in many manipulation tasks [1]. These strategies are applied to adapt grasping classifiers [18, 19, 20, 21, 22], disentangle object geometry without visual input [23, 24], and obtain object articulation constraints [25, 26]. Bauza and Rodriguez [27] develop an alternate method to compute uncertainty of push outcomes based on uncertainty of physical parameters. Many of these methods tackle information gathering by developing problem-specific solutions [21, 22], or relying on bandit-style exploration approaches [19, 20] or information-gain heuristics [18, 24] for short, single-step estimation tasks. Our work leverages POMDP solvers to tackle the information-gathering problem when multiple actions must be composed together.

### v3

@@ Abstract
Our goal is to enable robots to plan sequences of tabletop actions to push a block with unknown physical properties to a desired goal pose. We approach this problem by learning the constituent models of a Partially-Observable Markov Decision Process (POMDP), where the robot can observe the outcome of a push, but the physical properties of the block that govern the dynamics remain unknown. A common solution approach is to train an observation model in a supervised fashion, and do inference with a general inference technique such as particle filters. However, supervised training requires knowledge of the relevant physical properties that determine the problem dynamics, which we do not assume to be known. Planning also requires simulating many belief updates, which becomes expensive when using particle filters to represent the belief. We propose to learn an Attentive Neural Process that computes the belief over a learned latent representation of the relevant physical properties given a history of actions. To address the pushing planning problem, we integrate a trained Neural Process with a double-progressive widening sampling strategy. Simulation results indicate that Neural Process Tree with Double Progressive Widening (NPT-DPW) generates better-performing plans faster than traditional particle-filter methods that use a supervised-trained observation model, even in complex pushing scenarios.
@@ I Introduction
In scenarios where robots operate in unstructured, dynamic environments, planning under uncertainty is paramount. Some relevant properties for manipulation, such as object geometry, can be easily observed through sensors. Many other properties, such as inertial properties and friction, require interaction with the object [1]. The robot may not even know what the relevant physical properties are to determine action outcomes. One way robots can learn these inertial properties is to iteratively interact with objects and observe their behavior.
@@
We focus on an instance of this problem where a robot must compose a sequence of pushes to guide a block from a starting configuration to a goal configuration, without prior knowledge of the center of mass. Our goal is to develop a technique capable of inferring the true center of mass while simultaneously predicting outcomes of pushing actions.
@@
We model our planning challenge as a Partially Observable Markov Decision Process (POMDP), where unobservable physical properties remain constant over time [2]. Many POMDP solutions rely on Bayesian state estimation techniques (e.g., Kalman or particle filters) to update their beliefs as new observations become available. However, the linear assumptions required for Kalman filters do not hold in this scenario, since varying the center of mass can nonlinearly influence rotational outcomes. More general methods, such as particle filters, must be used, despite their high computational cost [3].
@@
To leverage these particle-based belief representations to solve our planning problem, a common method is to use Particle Filter Trees with Double Progressive Widening (PFT-DPW), a Monte Carlo tree search algorithm that performs a randomized search to identify the plan yielding the highest expected reward. In POMDP planning, it is crucial to balance exploration and exploitation [4]. PFT-DPW and similar Monte Carlo tree search methods manage this trade-off effectively by systematically exploring uncertain aspects of the system while exploiting known information to achieve high rewards [5].
However, PFT-DPW also has significant limitations due to the use of a particle filter. First, training an observation model used to update the particle filter requires knowing what the relevant physical properties of the object for the task dynamics are. Second, the runtime cost of updating a particle belief scales linearly with the number of particles when searching for long-horizon plans.
@@
To overcome these problems, we propose a model called the Pushing Neural Process (PNP), which jointly trains an inference network that estimates physical parameters, and a prediction network that learns to output a distribution of push outcomes based on physical properties. The inference network and prediction network are set up in an encoder-decoder architecture, where a representation of the physical properties is learned in the intermediate latent space. Notably, the PNP does not require any information about which physical properties are relevant to the problem, but instead encodes any relevant information about the representation into a multivariate gaussian outputted by the encoder.
@@
To compare the performance of both models, we derive a new planner that integrates the PNP, called NPT-DPW (Neural Process Tree with Double Progressive Widening). NPT-DPW replaces the particle filter used to maintain belief in PFT-DPW with a neural network that consumes the history of a given search and learns a representation for the object’s unknown physical properties. Our results show that, on average, NPT-DPW reaches up to twice as close to the goal location as PFT-DPW and avoids catastrophic failures (such as pushing into prohibited zones) far more often. We additionally demonstrate that the PNP’s latent space embeds a representation of the true COM by showing that the model becomes more accurate with greater amounts of previous pushes.
@@
Figure 1: The Push Neural Process architecture and its use as a subroutine within NPT-DPW. On the right of the image, the search tree over actions and outcomes is pictured, where at each step an action is selected. We then pass in the history of outcomes ($h$) that we have already observed into the encoder ($q_{\phi}$) of the model to obtain the distribution over the latent parameters, which we sample to get ($z$). We then pass $z$ to the decoder with our new action $a$ to get a distribution over outcomes. We sample from that distribution as an outcome ($o$) for the action $a$. We then add that outcome to the potential outcomes in the search tree.
@@ II Problem Setting
For the objects we aim to push, we categorize their properties into two distinct sets: observable properties, denoted by $x\in\mathcal{X}\subset\mathbb{R}^{n}$, and latent physical properties, denoted by $z\in\mathcal{Z}\subset\mathbb{R}^{m}$, where $m$ is decided empirically based on what results in the strongest model. The observable properties $x$ include features such as the block’s resting orientation, while $z$ captures physical properties such as center of mass. These latent properties are unknown to the robot a priori but play a critical role in determining the outcome of a push. We define a push in terms of an action $a\in\mathcal{A}$, characterized by an approach angle and a push velocity, where the angle is in the range $[0,2\pi)$. Upon completing a push, the robot observes the resulting pose of the block $o\in SE(2)$. A schematic of the push action is depicted in Figure 2.
Figure 2: A depiction of a push action $a\in\mathcal{A}$, parameterized by push angle $\theta\in[0,2\pi)$ and velocity $\overrightarrow{v}\in\mathbb{R}^{2}$. The center of mass $z$ (denoted as $\oplus$) is not directly observed.
@@ III-A Pushing Neural Process
The encoder consumes $D_{t}$ (of arbitrary length) along with the observable properties $x$ (object geometry) and infers a probability distribution over the latent parameters $z$. Although the latent space is uninterpretable, we hypothesize that the model will implicitly learn what useful information can be learned from pushing. To this end, we train a neural network, parameterized by $\phi$, to produce an approximate posterior $q_{\phi}(z|x,D_{t})$ that closely matches the true posterior $p(z|x,D_{t})$.
@@ III-B NPT-DPW
Instead of using a particle filter, our planner, NPT-DPW, uses the Pushing Neural Process to infer the underlying physical parameters and predict outcomes of future actions. To expand the tree, we pass the history of actions and outcomes through the encoder to infer the physical parameters. This history of actions implicitly represents a belief of the true properties, as the history of actions can be passed into the encoder to generate a distribution of the true physical properties.
Actions are sampled as they normally would be in PFT-DPW. Observations are sampled from the distribution computed by the decoder from the inferred parameters and the push under consideration. Thus, every step of the tree search will have runtime complexity of $O(h^{2})$, where $h$ is the size of the history (due to the quadratic complexity of self-attention), rather than $O(n)$, where $n$ is the number of particles used in PFT-DPW. We hypothesize a significant improvement in plan quality within a fixed computational budget, since $h\ll n$ in practice. As stated previously, an additional advantage of this method is that it works in general scenarios where the relevant properties are not known a priori to the robot.
@@ IV Experiments
Figure 4: This figure demonstrates the effectiveness of the NPT-DPW versus PFT-DPW (with 10/30/100 particles). Additionally, we conducted trials with 0.5,1,2,3,5, and 10 seconds of time to plan per step. The results plot the average percentage traveled towards the goal as a function of planning time allocated per action. For example, if the block started 3 meters away, and was pushed to being 1 meter way, that would be considered 66% progress. Fallen blocks are marked with the percentage distance corresponding to when they fell off. Therefore, the percentage from the goal varies a lot, since times the block will fall off almost immediately will be marked with an approximately 0% success.
@@
IV-A Scenarios
The experiments will be run using three tables with geometry designed to vary the hardness of the pushing problem (Figure 3). Each of these tables defines a different planning problem. The first is a blank-surface table designed to check whether the planner can produce plans that drive the block to the goal. In the second and third scenarios, the movement of the block must be constrained along certain paths to prevent it from falling off the pushing surface. The second table requires pushing in a straight line, while the third requires a push trajectory along a curve.
Solving these instances requires the planner to accurately infer the object’s physical properties. For the second and third tables, an intelligent planner must sequence information-gathering actions on a wider area before pushing the block along the narrow sections of these tables.
After each push, the robot obtains a reward for the block’s location $(x,y)$ relative to the goal $(g_{x},g_{y})$. We define the reward function
@@ IV-A Scenarios
fallen is a function that returns 1 if the block is off the table, and 0 otherwise. inRange is a function that is $1$ if the block is within 0.2 meters of the goal, and 00 otherwise. We set the discount factor $\gamma=0.8$ to incentivize the planner to prioritize not push the block off the table. An example of a plan generated by NPT-DPW is depicted in Fig. 5. While the metric used as the model to plan will be the reward function, we do not evaluate performance based on that, but instead use percent progress towards the goal, as can be seen later in the results section.
@@
We assess performance by fixing the computation time $t$ allowed at each step of the plan. Then for each $t$ in $[0.5,1,2,3,5,10]$ seconds, we ran 15 trials on each map for each planner, where we recorded the distance of the object from the goal after 30 steps. If the object fell off the table, or reached the goal early, we stopped then. We sampled a new center of mass for the block in each trial. For each tree search, we looked for the best action for depth equaling 3, and we only stored the first 10 actions for NPT-DPW, since empirically that was enough to gain a strong belief over the block.
@@
We also varied the number of particles in PFT-DPW, initializing it with $n=10,30$, or $100$, to evaluate NPT-DPW against a range of particle filter–based planners and assess the impact of particle count on performance. Additionally, while maintaining the particle filter, if the weight associated with a particle became smaller than $10^{-9}$, then we turn off that particle, which should encourage the particle filter to become faster as plans become deeper, since low weighted particles will no longer be counted.
@@
Figure 5: This is an example of a plan created by NPT-DPW and run by the simulation on the ”ring” map. Important features that show signs of intelligent planning are the pushes in the beginning in the corner, signifying learning more about the block before trying to move towards the goal. Note the figure has been re-scaled slightly from the real experiments to be more visually clear.
These simulations are run in the PyBullet simulator. We implement an end-effector PD-controller to ensure the gripper’s end-effector motion velocity is consistent throughout the motion. To sample actions for progressive widening, we uniformly sample an angle from $[0,2\pi]$ in world coordinates and push at the velocity of $0.10$ m/sec. Each of the blocks is a rectangular prism with a fixed geometry of 2.5 cm x 2.5 cm x 1.5 cm. The mass of the block is set constant at 0.2 kg. The center of mass of the block is sampled uniformly within the convex hull of the object’s geometry.
Our neural network architecture is based on Attentive Neural Processes [6]. Each push is first encoded by an MLP and processed through multiple layers of self-attention. After that, the outputs of the self-attention layers are aggregated into outputs consisting of a mean and covariance. These represent the mean and diagonal covariance of a multivariate Gaussian over $z$, representing the distribution over latent physical properties. Note that for our problem, the only latent physical properties are the $x$ and $y$ coordinates of the center of mass, but we still set $z\in\mathbb{R}^{5}$ since it empirically led to improved performance. The decoder is structured as multiple linear layers that consume a sample from the latent space and outputs a distribution over possible outcome poses of the block after the push is finished.
TABLE I: Number of Simulations Done with Different Amounts of Planning Time, Averaged Over 10 Plans
Planner 1 s 5 s 10 s
NPTDPW $325.3$ $2291$ $4807$
PFTDPW with 10 Particles $157.6$ $806.6$ $1910$
PFTDPW with 30 Particles $62.80$ $646.0$ $1410$
PFTDPW with 100 Particles $30.20$ $322.2$ $977.1$
@@ IV-B Results
Our results, shown in Figure 4, highlight the effectiveness of using a learned latent space for efficient planning. NPT-DPW consistently outperforms PFT-DPW across multiple scenarios when allocated at least one second per action. This performance gain suggests that the latent space learned by the Attentive Neural Process provides a compact yet expressive understanding of the block’s unknown physical properties, which NPT-DPW was able to leverage to achieve greater success planning.
Figure 6: This plot shows how the average error of the predicted location of the block from the Push Neural Process to the true result of the location of that push decreases as the model receives more past interactions, or context pushes. The error bars represent an interval of one standard error around the mean distance.
Figure 6 demonstrates that knowing the result of previous interactions with a block (called ”context pushes”) enables the robot to make more accurate pushes, since the robot is able to process those previous outcomes in the encoder. Because of this, we can infer that the PNP’s latent space is able to learn a representation of the block’s center of mass.
In addition, maintaining belief through the latent space reduces the computational overhead of having runtime proportional to the number of particles of explicit particle representations, which is what we hypothesize allows NPT-DPW to more rapidly explore the space of potential action sequences than PFT-DPW. In specific, in Table I, we have shown that the number of planning iterations the PNP can process is significantly higher than the particle filter even when the number of particles is 10.
While all these results are extremely promising, there are a few caveats. In scenarios where very limited planning time (e.g., half a second per action) was available, both models struggled similarly, underscoring the necessity for adequate computation time to fully leverage the benefits of learned latent spaces. NPT-DPW, however, starts out performing PFT-DPW when the amount of computation becomes greater than a second.
Another interesting result of the data is the lack of difference in performance between the varied number of particles in PFT-DPW. We hypothesize this is due to the nature of the pushing task, where subtle differences in the center of mass have limited impact because push outcomes primarily depend on the relative positions of the center of mass and the point of contact. As a result, a small number of particles might already capture most of the informative uncertainty, while increasing particle counts does not significantly enhance prediction quality but rather increases computational load, limiting the planner’s capacity for deeper exploration.
Our findings reinforce the potential of latent spaces in the Attentive Neural Process to maintain a useful learned representation of the unobservable properties of an object. Additionally, we showed that using the Attentive Neural Process in a planner leads to significantly improved performance compared to PFT-DPW. These results indicate promising directions for future exploration into richer representation learning, aiming to enhance generalization and robustness in robotic planning tasks.
@@ V Related Work
Estimating unobserved properties through interaction is a common strategy in many manipulation tasks [1]. These strategies are applied to adapt grasping classifiers [18, 19, 20, 21, 22], disentangle object geometry without visual input [23, 24], and obtain object articulation constraints [25, 26]. Bauza and Rodriguez [27] develop an alternate method to compute uncertainty of push outcomes based on uncertainty of physical parameters of known relevance. Many of these methods tackle information gathering by developing problem-specific solutions [21, 22], or relying on bandit-style exploration approaches [19, 20] or information-gain heuristics [18, 24] for short, single-step estimation tasks. Our work leverages POMDP solvers to tackle the information-gathering problem when multiple actions must be composed together.